OpenStudy (anonymous):

will give medals which equation represents the axis of symmetry of the function y = -2x^2 + 4x - 6 A. y = 1 B. x = 1 C. x = 3 D. x = -3

OpenStudy (anonymous):

First off, what is the axis of symmetry?

OpenStudy (anonymous):

I don't know

OpenStudy (anonymous):

Well then. The axis of symmetry is an x value that points to the middle of the equation.

OpenStudy (anonymous):

In other words, it is the x value of your parobola vertex. Find the vertex of your parabola.

OpenStudy (anonymous):

The axis of symmetry can be found using: \[x=-\frac{ b }{ 2a }\] This works for any quadratic equation of the form: \[y = ax ^{2}+bx+c\] In your problem: a = -2 b = 4 and c = -6

OpenStudy (anonymous):

\[x = - \frac{ b }{ 2a }= -\frac{ 4 }{ 2(-2) }=-\frac{ 4 }{ -4 }=1\]

OpenStudy (anonymous):

so the answer is x = 1?

OpenStudy (anonymous):

Yes!!

OpenStudy (anonymous):

x=1

OpenStudy (anonymous):

thank you to everyone who helped